Sol Summary – February 07th

Crew 240 Sol Summary Report 07Feb2022

Sol: 7

Summary Title: Getting back in gear

Author’s name: Clément Plagne

Mission Status: Nominal

Sol Activity Summary: Following a week spent toying with the original mission planning to adjust to varying EVA times and duration, and a weekend of rest breaking the rhythm, the most important part of starting this second of three weeks is getting a strong rhythm back and sticking to the plan. With a day heavy with an EVA, several human sciences experiments to perform, 3D printing to start, and a lot more, timekeeping was of the essence.

Fortunately for us, everything went smoothly – the EVA went without an issue (and I’ve got to add, seeing crew members up on Hab Ridge from the inside is always fun), there have been massive upswings for the work of Maxime, our astronomer, who is now busier than ever, and we still managed to fit our zillion activities well enough so that our HSO could have us get a small meditation session before heading out to write reports.

Look Ahead Plan: A slightly less busy day tomorrow, with some leeway to have some more relaxing activities during the afternoon. Another crack at the North Ridge with Pierre’s updated and well processed 3D data. Finally fixing the missing piece of one of our atmospheric study experiments so we can take it out in the field at last.

Anomalies in work: None to report

Weather: Fair and warm.

Crew Physical Status: crewmember who had hurt themselves in a fall are doing better, with some bruises

EVA: One performed, went smoothly.

Reports to be filed: Sol Summary, Astronomy Report, Operations Report, EVA Report, EVA Request, Journalist Report, Crew Pictures, HSO Report

Support Requested: None in particular. Thanks for the salt!

Sol Summary – Feb 6th

Crew 240 Sol Summary Report 06Feb2022

Sol: 6

Summary Title: A well deserved rest

Author’s name: Clément Plagne

Mission Status: Nominal

Sol Activity Summary: All of us slept in except myself, who woke up at 5am to take advantage of the night’s extra data to download data for Maxime’s remote astronomy observations. Fortunately for me, I went right back to sleep for a couple of hours. Unfortunately for Maxime, the literal gigabytes of data we’d accumulated were of no use, as the telescope had issues and only recorded noise. It was with great luck that he at least managed to catch a few images of the Sun during the day while the Musk telescope was pointed the right way.

Days are seldom fully idle however, and we still got a small routine EVA done for battery change, finally achieved or first 3D printed objects, and a part of the crew spent the afternoon doing questionnaires and tests for the work we do with two French universities. Nevertheless, it was still overall a fairly relaxing day, and we’re looking forward to pizza night for dinner.

Look Ahead Plan: Pizza night for dinner. (also, using the 3D printed piece for the water recycling and printing new pieces that we had in our backlog). Scheduling some remote astronomy observations that will hopefully function well.

Anomalies in work: Issues with the remote observations made at the observatory in New Mexico – of 6 observations made, none of them had functional data, possibly owing to an issue with the dome or the sensors.

Weather: Fair.

Crew Physical Status: A crewmember had a fall in the stairs between lower and upper deck, and got away with some bruises and pain but nothing major, very fortunately.

EVA: One, very efficient, only to switch batteries on an atmospheric science instrument.

Reports to be filed: Astronomy report, Science Report, EVA Report, EVA request, Operations Report, HSO Report, Sol Summary, Journalist Report, Crew Photos

Support Requested: None

Sol Summary – Feb 4th

Crew 240 Sol Summary Report 05Feb2022

Sol:5

Summary Title:

Author’s name: Clément Plagne

Mission Status: Nominal

Sol Activity Summary: Since our workouts are only on weekdays and the EVA was a little late in the morning, most of the crew enjoyed sleeping in – to some degree, as most awoke by 8:30, and I somehow stuck to my usual 6:45 by what I assume was sheer force of habit. I also found time to get some pancakes going – thanks Shannon for the advice, they were perfect!

The EVA was one I’d done two years ago: stepping into Candor Chasma. I’d always thought the term “unforgettable” was fuzzy, and more figurative than anything else. This proved me wrong. I remembered rocks I’d stepped on, trails I’d walked on and formations I’d photographed years ago like it was yesterday. I’m happy the rest of the EVA team found it as remarkable as I did.

This start of the weekend after this tiring week, a big and the occasion to sleep in meant that most of the hab was idling in bed for the better part of the afternoon. Nevertheless, recycling of shower water chugged along well, and the 3D printer finally ran correctly – I’d diagnosed an issue with the mechanics of motion of the system along the Y axis, and Maxime and François cleverly fixed it. We’re planning an evening based on cohesion games and relaxation exercises.

Look Ahead Plan: A good night of sleep tomorrow morning. Started 3D printing of a piece I need to setup an experiment outside.

Anomalies in work: None noticed

Weather: Fair, end of EVA felt quite hot inside the suits but not surprisingly so.

Crew Physical Status: The extra sleeping time and rest really improved physical and mental states across the crew.

EVA: One to Candor Chasma, which went without any significant issue besides us noticing some human activity along the way and taking steps to evade it.

Reports to be filed: Sol Summary, Operations Report, HSO Report, Journalist Report, Crew Photos, EVA Report, EVA Request, Astronomy Report

Support Requested: None

Sol Summary – Feb 4th

Crew 240 Sol Summary Report 04Feb2022

Sol: 4

Summary Title: Ups and Downs

Author’s name: Clément Plagne

Mission Status: All nominal

Sol Activity Summary:

Day started by getting the toilet back in good working order as given by Shannon, with no issue in particular.

After looking at it from below and waiting for Pierre to 3D scan it, we finally attempted a climb of North Ridge today. Doing this from the North, to give us an unfamiliar start, it took many attempts find a suitably safe place to begin the ascent, which unfortunately was cut short after I dropped my radio and lost contact with the rest of the team. Return was fine, battery change on the experiments went well, and in all, the EVA was still quite successful and honestly a solid workout. As a bonus, the crewmember that had bad back pain is now feeling great, thanks to some adjustments to the suit – thanks Scott!

We think we’ve finally go to the bottom of the 3D printer issue, and we now think it’s mostly a mechanical thing caused by both the hits the printer must have taken during its flight a year back, and the low temps in the hab. We’ll be trying to fix it. In the meantime, Maxime got his first solar observations today – the dome rotation is limited (a few degrees before locking up), and we think something might be up with the gears (battery wasn’t the issue), but observations could be done at the right hour. Julie also started up on some botany in the GreenHab.

Look Ahead Plan: For the third day in a row, I know, but getting something out of the 3D printer. Trying to put out a good exploration/lookout EVA, and getting a smidge more rest as we’re not doing the morning workouts on the weekends.

Anomalies in work: None to report.

Weather: Almost warm during the EVA but comfortable

Crew Physical Status: Fine. Pierre had a small nosebleed during the EVA, but barely noticed it.

EVA: One done in the afternoon, partial success (return home before reaching final point due to user discomfort). Issues with headsets during EVA.

Reports to be filed: Sol Summary, Crew Photos, Journalist Report, HSO Report, Ops Report, EVA Report, EVA Request (note: is an Astronomy report warranted if Maxime just used the observatory?)

Support Requested: We’ve been trying to cook basic pancakes on the skillet for days now and it’s been hell all the way through – no matter how much oil or whatever the heat, it sticks like hell. This might seem trivial but I still feel it’s a shame not to be able to do this sort of basic cooking. Are there any other cooking implements available?

Sol Summary – Feb 3th

Crew 240 Sol Summary Report 03Feb2022

Sol: 3

Summary Title: A little frustration

Author’s name: Clément Plagne

Mission Status: All nominal

Sol Activity Summary: Having moved the planned EVA to the afternoon to avoid the very cold temperatures of the morning, we set out to work on our experiments over the morning. The 3D printer experiment still isn’t running well, probably because the printer had been in storage for close than a year prior to the mission. At the same time, Maxime’s Astronomer trouble continued – with the Robotic Observatory in reparations and the computer used for solar observations being broken, it was already a struggle to get some work in. Unfortunately, there were further issues in the Musk Observatory, and Maxime’s astronomy work was hindered even more.

During EVA, the 3D mapping work started by Pierre was successfully completed. However, issues with headsets and significant discomfort with the suits led to the EVA being cut short before the secondary objective of climbing up Hab Ridge could be accomplished.

There’s still a while to go, we’ll get things running.

Look Ahead Plan: Getting a good look at Pierre’s dynamic map and further processing that data, getting that darn 3D printer working

Anomalies in work: Issues with the Musk Observatory noted.

Weather: Very cold morning, but comfortable, or even warm afternoon for the EVA team

Crew Physical Status: Fine. Some nosebleeds, and a little EVA fatigue

EVA: One done in the afternoon, partial success (return home before reaching final point due to user discomfort). Issues with headsets during EVA.

Reports to be filed: Sol Summary, Astronomy Report, Crew Photos, Journalist Report, HSO Report, Ops Report, EVA Report, EVA Request

Support Requested: If there’s any possibility to have that, replacements to the headsets could be useful

Sol Summary – February 2nd

Crew 240 Sol Summary Report 02Feb2022

Sol: 2

Summary Title: Always walk before you run.

Author’s name: Clément Plagne

Mission Status: Business as usual, getting into the rhythm

Sol Activity Summary: The morning was dedicated to setting up two out of three pieces of our atmospheric equipment: the LOAC (an aerosol particle counter) and its associated weather station – our electric field instrument had a piece of it misplaced on our way there, and will need a little handiwork to get back into a usable state. Installation was slowed down by the fact that the weather station had many parts of its setup that we didn’t handle as well as we wished, and it’ll need touching upon a further EVA.

Inside the Hab, Julie has been working on a way to clean up some of the water we use, so we can reuse some of it on non-crucial things. This eventually required us to custom print some pieces with our 3D printer. Getting that one working well was another tedious task, which we barely accomplished by the end of the day.

In the meantime, Pierre finalized data analysis on his 3D map of the North Ridge and concluded that he needed another go at it to get more usable data. On the good news side, our Engineer fixed up my spacesuit handily.

In all, things have had some trouble getting going. This is as expected for this sort of mission – this is a big, clunky machine we’re getting going! We’ll go on and improve as we go along with it.

Look Ahead Plan: Getting our first 3D printed parts out, getting a better 3D map of the Hab ridge, and getting some data from the atmospheric experiments.

Anomalies in work: We mistakenly plugged our rovers back in after today’s EVA, and didn’t check the SOC frequently enough, leading to the level getting abnormally low, that Atila fixed. We’ve modified procedures and set up new ones to ensure this doesn’t occur again.

Weather: Cloudy, but felt warm on EVA.

Crew Physical Status: All is well

EVA: Occurred without any anomalies, was a near-total success

Reports to be filed: Sol Summary, HSO Report, Ops Report, Journalist Report, EVA Report, EVA Request, Pictures

Support Requested: None

Sol Summary – January 31st

Sol Activity Summary: Morning was spent with Atila, finalizing the steps required for us to safely and confidently perform EVAs. Following a very substantial lunch, we took our final pictures outside and entered sim mid-afternoon. During the following time, we continued setting up various science experiments and began getting used to life in sim.

Look Ahead Plan: Planning for and performing our first EVAs in order to get our footing outside.

Anomalies in work: Nothing of note.

Weather: Slightly overcast, biting cold in the morning.

Crew Physical Status: No health issues, a little fatigue for some

EVA: None performed besides outside training with Atila

Reports to be filed: Operations Report, Health and Safety Officer’s report (with temperatures), Pictures, Journalist’s report, Sol Summary, EVA requests (2 if possible)

Support Requested: We have taken pictures of the visible insulation mentioned previously and can supply them if need be.

Sol Summary – Jan 28th

Sol: 11

Summary Title: Goodbye Mars

Author’s name: Yael Méndez- Commander

Mission Status: The entire crew is happy to return to Earth, however, we will all miss living on Mars.

Sol Activity Summary: The crew got up early to prepare the station for the next rotation. We had pizza and salad for lunch to celebrate the end of our simulation.

Look Ahead Plan: Back home.

Anomalies in work: None.

Weather: Sunny in the morning and afternoon.

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, Operations Report, Daily photos, Journalist Report, Crew 226 Food and Supply Inventory, Mission Summary Crew 226

Support Requested: None

Sol Summary

Sol: 10

Summary Title: Mars, the inhospitable red planet

Author’s name: Yael Méndez- Commander

Mission Status: A day to carry out tasks on the station and to continue learning about the Geology of Mars. After a power failure, it got restored at 7 pm.

Sol Activity Summary: We started the day with a breakfast of oatmeal with fruits. We visited again Phobos Peak to take some samples of clays and sands. We experienced a power failure on Mars.

Look Ahead Plan: End SIM and preparing our trip back to Earth.

Anomalies in work: SOC Power failure.

Weather: Sunny and Windy in the morning and afternoon.

Crew Physical Status: Nominal

EVA: One EVA realized

Reports to be filed: Sol Summary, Operations Report, EVA Report, HSO Report, Daily photos, Journalist report

Support Requested: None

Sol Summary – January 26th

Sol: 9

Summary Title: Exploring Mars 🔴

Author’s name: Yael Méndez- Commander

Mission Status: A day to carry out tasks on the station and to continue learning about the Geology of Mars.

Sol Activity Summary: We started the day with a tasty egg omelet breakfast.

We visited Phobos Peak to learn about the soil conditions of the area and take some samples of clays and sands.

Look Ahead Plan: Continue working on mission projects and preparing our trip back to Earth.

Anomalies in work: None

Weather: Sunny in the morning and afternoon.

Crew Physical Status: Nominal

EVA: One EVA realized

Reports to be filed: Sol Summary, Operations Report, EVA Report, HSO Report, Daily photos, Journalist report

Support Requested: Dehydrated lime